# CultureGuard: Towards Culturally-Aware Dataset and Guard Model for Multilingual Safety Applications

### v3

@@ 1 Introduction
The advancements in Large Language Models (LLMs) have led to their widespread adoption in diverse real-world agentic applications. Both closed-source models, such as ChatGPT Achiam et al. (2023), Gemini Team et al. (2023), and Claude, and their rapidly evolving open-source counterparts, including Nemotron Adler et al. (2024), Llama Dubey et al. (2024), Gemma Team et al. (2025), Mistral Jiang et al. (2024), and Qwen Yang et al. (2025), have demonstrated significant capabilities in solving complex problems. As the applications of LLMs continue to expand rapidly, the need for these models to be developed and deployed responsibly and safely has grown proportionally.
@@
Ensuring LLM safety typically involves two primary strategies: integrating refusal samples into the model’s alignment data during training, and implementing post-deployment auditing of LLM inputs and outputs via specialized Safety Guard models Wang et al. (2025a). This work specifically focuses on the latter, with a particular emphasis on multilingual safety guard models. It has been critically observed that LLMs demonstrate an elevated tendency to produce unsafe content in languages other than English Wang et al. (2024). However, the development of robust multilingual guard models is severely hampered by significant resource constraints in non-English linguistic contexts. The creation of culturally nuanced safety datasets, which necessitates expert annotators from each specific region, renders the data collection process costly. Consequently, research and development in multilingual LLM safety guards remain limited.
Early endeavors in this domain primarily concentrated on multilingual safety guard evaluation datasets Deng et al. ; de Wynter et al. (2025). While models like LlamaGuard Chi et al. (2024), ShieldGemma Zeng et al. (2024), and Granite Guardian Padhi et al. (2024) offer some support for non-English languages, their performance is often suboptimal compared to their English counterparts. Furthermore, the proprietary nature of their datasets and methodologies restricts their utility and customizability for broader research and application. Similarly, open models like Llama-Nemotron-Safety-Guard-Defensive-V1 Ghosh et al. (2024), Llama-Nemotron-Safety-Guard-V2 Ghosh et al. (2025), and WildGuard Han et al. (2024) also lack dedicated multilingual support. More recent contributions, such as PolyGuard Kumar et al. (2025), DuoGuard Deng et al. (2025), and OmniGuard Verma et al. (2025), have emerged. While PolyGuard is a valuable contribution to multilingual safety, its dependence on labels generated by GPT-4o could raise issues regarding its commercial viability. OmniGuard, on the other hand, is a model-specific solution that presents practical challenges for broader applicability. DuoGuard supports only four languages and does not provide consistent performance across benchmarks. Moreover, these models predominantly leverage machine translation without adequately addressing the crucial aspect of cultural relevance, which is paramount for effective cross-cultural safety.
@@
In response to these challenges, we present CultureGuard, a novel framework designed for curating culturally aligned, high-quality safety datasets across multiple languages. This method is purely based on synthetic data generation, leveraging the English generation capabilities of LLMs and strong translation methods. Importantly, it does not require human input, which presents a scalable approach for targeting multiple languages or regions. Our technical approach employs a four-stage pipeline, encompassing cultural data segregation, cultural data adaptation, machine translation, and quality filtering. Cultural adaptation of the samples using LLMs to the target culture is central to the CultureGuard approach; its core components are shown in Figure 2. It is important to note that since this adaptation is performed on the English source text prior to translation, our approach focuses on ideational and social cultural elements, while linguistic aspects are not explicitly covered Liu et al. (2025). We also present a novel cross-lingual consistency filter to remove low-quality samples post-translation. This pipeline facilitates the systematic conversion and expansion of the English Nemotron-Content-Safety-Dataset-V2 (formerly known as Aegis 2.0) safety dataset Ghosh et al. (2025) into eight distinct languages: Arabic, German, Spanish, French, Hindi, Japanese, Thai, and Chinese (Mandarin). Additionally, we synthetically curate jail-breaking prompt-response pairs using LLMs to enhance the detection capabilities for such adversarial inputs. Safety labels for these synthetic pairs are computed via a jury of LLMs. The final dataset, termed Nemotron-Safety-Guard-Dataset-v3111Nemotron-Safety-Guard-Dataset-v3, consists of a total of 386,661 samples across 9 languages (including English) as shown in Figure 3. The dataset is based on the English Nemotron-Content-Safety-Dataset-V2 and uses the same safety risk taxonomy.
@@ 2 Related Work
Recent progress in multilingual content safety has centered around the development of comprehensive datasets and robust evaluation methodologies. Key benchmarks include XSafety Wang et al. (2024), MultiJail Deng et al. , RTP-LX de Wynter et al. (2025), PolygloToxicityPrompts (PTP) Jain et al. (2024) and the Aya Red-teaming Ahmadian et al. (2024) datasets. These resources typically combine naturally occurring human–LLM interactions with human-verified machine translations to provide a more holistic assessment of LLM safety capabilities beyond English. They evaluate a range of dimensions, from general harmfulness and toxicity to jailbreak susceptibility and culturally specific safety violations.
Parallel advancements in multilingual guard models have been instrumental in addressing content moderation challenges across languages. PolyGuard Kumar et al. (2025) emphasizes the creation of large, diverse multilingual datasets using machine translation augmented with human validation, which are then used to train dedicated safety moderation models. DuoGuard Deng et al. (2025) addresses data scarcity through a two-player reinforcement learning setup, enabling the adversarial generation of high-quality synthetic multilingual safety data. OmniGuard Verma et al. (2025) focuses on efficiency and modality-awareness by leveraging the cross-lingual and cross-modal internal representations of LLMs/MLLMs to construct lightweight harmfulness classifiers that require minimal inference cost.
Recent work has also explored smaller, more task-adaptive safety models. For example, STAND-Guard Wang et al. (2025b) demonstrates competitive multilingual performance despite a smaller parameter footprint, suggesting the feasibility of compact yet effective safety solutions. Similarly, SELF-DEFENCE Deng et al. presents a promising framework for reducing unsafe generation rates by automatically generating multilingual safety training data for fine-tuning. This approach proves effective in mitigating both unintentional and adversarial harms.
Despite these advances, significant challenges remain. Notably, issues such as data imbalance, the "harmfulness curse" in low-resource languages, and the cross-lingual leakage of unsafe behaviors within multilingual LLMs Lu and Koehn (2024); Huang et al. (2024) continue to impede progress. Moreover, existing models rely heavily on translated datasets without ensuring cultural fidelity. These limitations underscore the critical need for culturally aware, scalable solutions in multilingual safety research, a gap that our work, CultureGuard, aims to address.
@@ 3.1.1 Cultural Data Segregation
We first identify and segregate culturally influenced samples from generic samples within the Nemotron-Content-Safety-Dataset-V2 English dataset. Culturally influenced data is specifically tied to a particular region, country, or cultural group. This segregation is achieved by prompting the Llama-3.1-Nemotron-70B-Instruct LLM to classify text based on whether it includes references, traditions, language, or practices unique to any specific region or culture, distinct from general content that lacks such ties. For example, "Any other Jew joke you got?" would be considered a cultural sample as its relevance may vary across different target cultures (e.g., it may not be relevant in an Indian context). Conversely, "What steps should I take to rob a house?" is classified as a generic sample, as such content is universally relevant across regions. The classification focuses on regional or cultural specificity, without considering tone, appropriateness, offensiveness, or politeness of the content. To validate the accuracy of this segregation, we utilized GPT-4o as an independent judge. Our cultural segregation method achieved a test accuracy of 95.32% when using GPT-4o’s labels as ground truth.
@@ 3.1.2 Cultural Data Adaptation
Cultural adaptation is critical for content safety as perceptions of harmfulness are deeply rooted in cultural context. Segregated culturally specific English data undergoes adaptation to target linguistic and cultural domains. This adaptation is carried out on English text as existing open LLMs do not perform very well on non-English text. This process is facilitated by Mixtral-8x22B, which acts as a culturally aware editor, ensuring content authenticity and relevance within the target cultural context. Mixtral-8x22B is chosen because it does not refuse to adapt harmful samples and does not have any license restrictions. The LLM is tasked with adapting the content to make it relevant and relatable to a specific target culture (e.g., Indian region and culture), ensuring the original theme, message, and intent remain unchanged. This involves modifying names, examples, locations, idioms, festivals, food, clothing, traditions, and other culturally specific references to reflect the target cultural contexts. For instance, "Any other Jew joke you got?" might be adapted to "Any other Sardar joke you got?" for an Indic context. Crucially, the process ensures that the adapted content is naturally localized and culturally appropriate, without altering the inherent safety level or applying any safety judgments or ethical filters from the original content.
@@ 3.1.4 Quality Filtering
To mitigate the second problem of translation mistakes changing meaning, we employ FAITH-based filtering Paul et al. (2025). We use an LLM Judge to rate the translations based on five crucial aspects: Fluency, Accuracy, Idiomaticity, Terminology, and Handling of Format (FAITH). We use a low filtering threshold (3.5/5) here so that minimal samples are being filtered out.
@@
Table 1: Performance comparison of safety guard models on various multilingual safety benchmarks, measured by harmful-F1 score. CultureGuard model variations in this table refer to the Llama-3.1-Nemotron-Safety-Guard-8B-v3 model. CultureGuard (vanilla) utilizes a translation and filtering pipeline. CultureGuard additionally incorporates cultural adaptation. JB SDG refers to the use of multilingual jailbreak Synthetic Data Generation, and QF denotes quality filtering. Benchmark abbreviations: CultureGuard (Nemotron-Safety-Guard-Dataset-v3), PGPrompts (PolyGuardPrompts). Nemotron-Safety-Guard-V2 refers to Llama-Nemotron-Safety-Guard-V2.

### v4

@@ 1 Introduction
The advancements in Large Language Models (LLMs) have led to their widespread adoption in diverse real-world agentic applications. Both closed-source models, such as ChatGPT achiam2023gpt, Gemini team2023gemini, and Claude, and their rapidly evolving open-source counterparts, including Nemotron adler2024nemotron, Llama dubey2024llama, Gemma team2025gemma, Mistral jiang2024mixtral, and Qwen yang2025qwen3, have demonstrated significant capabilities in solving complex problems. As the applications of LLMs continue to expand rapidly, the need for these models to be developed and deployed responsibly and safely has grown proportionally.
@@
Ensuring LLM safety typically involves two primary strategies: integrating refusal samples into the model’s alignment data during training, and implementing post-deployment auditing of LLM inputs and outputs via specialized Safety Guard models wang2025survey. This work specifically focuses on the latter, with a particular emphasis on multilingual safety guard models. It has been critically observed that LLMs demonstrate an elevated tendency to produce unsafe content in languages other than English wang2024all. However, the development of robust multilingual guard models is severely hampered by significant resource constraints in non-English linguistic contexts. The creation of culturally nuanced safety datasets, which necessitates expert annotators from each specific region, renders the data collection process costly. Consequently, research and development in multilingual LLM safety guards remain limited.
Early endeavors in this domain primarily concentrated on multilingual safety guard evaluation datasets dengmultilingual; de2025rtp. While models like LlamaGuard chi2024llama, ShieldGemma zeng2024shieldgemma, and Granite Guardian padhi2024granite offer some support for non-English languages, their performance is often suboptimal compared to their English counterparts. Furthermore, the proprietary nature of their datasets and methodologies restricts their utility and customizability for broader research and application. Similarly, open models like Llama-Nemotron-Safety-Guard-Defensive-V1 ghosh2024aegis_1, Llama-Nemotron-Safety-Guard-V2 ghosh2025aegis2, and WildGuard han2024wildguard also lack dedicated multilingual support. More recent contributions, such as PolyGuard kumar2025polyguard, DuoGuard deng2025duoguard, and OmniGuard verma2025omniguard, have emerged. While PolyGuard is a valuable contribution to multilingual safety, its dependence on labels generated by GPT-4o could raise issues regarding its commercial viability. OmniGuard, on the other hand, is a model-specific solution that presents practical challenges for broader applicability. DuoGuard supports only four languages and does not provide consistent performance across benchmarks. Moreover, these models predominantly leverage machine translation without adequately addressing the crucial aspect of cultural relevance, which is paramount for effective cross-cultural safety.
@@
In response to these challenges, we present CultureGuard, a novel framework designed for curating culturally aligned, high-quality safety datasets across multiple languages. This method is purely based on synthetic data generation, leveraging the English generation capabilities of LLMs and strong translation methods. Importantly, it does not require human input, which presents a scalable approach for targeting multiple languages or regions. Our technical approach employs a four-stage pipeline, encompassing cultural data segregation, cultural data adaptation, machine translation, and quality filtering. Cultural adaptation of the samples using LLMs to the target culture is central to the CultureGuard approach; its core components are shown in Figure 2. It is important to note that since this adaptation is performed on the English source text prior to translation, our approach focuses on ideational and social cultural elements, while linguistic aspects are not explicitly covered liu2025culturally. We also present a novel cross-lingual consistency filter to remove low-quality samples post-translation. This pipeline facilitates the systematic conversion and expansion of the English Nemotron-Content-Safety-Dataset-V2 (formerly known as Aegis 2.0) safety dataset ghosh2025aegis2 into eight distinct languages: Arabic, German, Spanish, French, Hindi, Japanese, Thai, and Chinese (Mandarin). Additionally, we synthetically curate jail-breaking prompt-response pairs using LLMs to enhance the detection capabilities for such adversarial inputs. Safety labels for these synthetic pairs are computed via a jury of LLMs. The final dataset, termed Nemotron-Safety-Guard-Dataset-v3111Nemotron-Safety-Guard-Dataset-v3, consists of a total of 386,661 samples across 9 languages (including English) as shown in Figure 3. The dataset is based on the English Nemotron-Content-Safety-Dataset-V2 and uses the same safety risk taxonomy.
@@ 2 Related Work
Recent progress in multilingual content safety has centered around the development of comprehensive datasets and robust evaluation methodologies. Key benchmarks include XSafety wang2024all, MultiJail dengmultilingual, RTP-LX de2025rtp, PolygloToxicityPrompts (PTP) jain2024polyglotoxicityprompts and the Aya Red-teaming ahmadian2024multilingual datasets. These resources typically combine naturally occurring human–LLM interactions with human-verified machine translations to provide a more holistic assessment of LLM safety capabilities beyond English. They evaluate a range of dimensions, from general harmfulness and toxicity to jailbreak susceptibility and culturally specific safety violations.
Parallel advancements in multilingual guard models have been instrumental in addressing content moderation challenges across languages. PolyGuard kumar2025polyguard emphasizes the creation of large, diverse multilingual datasets using machine translation augmented with human validation, which are then used to train dedicated safety moderation models. DuoGuard deng2025duoguard addresses data scarcity through a two-player reinforcement learning setup, enabling the adversarial generation of high-quality synthetic multilingual safety data. OmniGuard verma2025omniguard focuses on efficiency and modality-awareness by leveraging the cross-lingual and cross-modal internal representations of LLMs/MLLMs to construct lightweight harmfulness classifiers that require minimal inference cost.
Recent work has also explored smaller, more task-adaptive safety models. For example, STAND-Guard wang2025stand demonstrates competitive multilingual performance despite a smaller parameter footprint, suggesting the feasibility of compact yet effective safety solutions. Similarly, SELF-DEFENCE dengmultilingual presents a promising framework for reducing unsafe generation rates by automatically generating multilingual safety training data for fine-tuning. This approach proves effective in mitigating both unintentional and adversarial harms.
Despite these advances, significant challenges remain. Notably, issues such as data imbalance, the "harmfulness curse" in low-resource languages, and the cross-lingual leakage of unsafe behaviors within multilingual LLMs lu2024learn; huang2024survey continue to impede progress. Moreover, existing models rely heavily on translated datasets without ensuring cultural fidelity. These limitations underscore the critical need for culturally aware, scalable solutions in multilingual safety research, a gap that our work, CultureGuard, aims to address.
@@ 3.1.1 Cultural Data Segregation
We first identify and segregate culturally influenced samples from generic samples within the Nemotron-Content-Safety-Dataset-V2 English dataset. Culturally influenced data is specifically tied to a particular region, country, or cultural group. This segregation is achieved by prompting the Llama-3.1-Nemotron-70B-Instruct LLM to classify text based on whether it includes references, traditions, language, or practices unique to any specific region or culture, distinct from general content that lacks such ties. For example, "How can I get a fake Social Security card to get a job?" would be considered a cultural sample as its relevance may vary across different target cultures/regions (e.g., it may not be relevant in an Indian context). Conversely, "What steps should I take to rob a house?" is classified as a generic sample, as such content is universally relevant across regions. The classification focuses on regional or cultural specificity, without considering tone, appropriateness, offensiveness, or politeness of the content. To validate the accuracy of this segregation, we utilized GPT-4o as an independent judge. Our cultural segregation method achieved a test accuracy of 95.32% when using GPT-4o’s labels as ground truth.
@@ 3.1.2 Cultural Data Adaptation
Cultural adaptation is critical for content safety as perceptions of harmfulness are deeply rooted in cultural context. Segregated culturally specific English data undergoes adaptation to target linguistic and cultural domains. This adaptation is carried out on English text as existing open LLMs do not perform very well on non-English text. This process is facilitated by Mixtral-8x22B, which acts as a culturally aware editor, ensuring content authenticity and relevance within the target cultural context. Mixtral-8x22B is chosen because it does not refuse to adapt harmful samples and does not have any license restrictions. The LLM is tasked with adapting the content to make it relevant and relatable to a specific target culture (e.g., Indian region and culture), ensuring the original theme, message, and intent remain unchanged. This involves modifying names, examples, locations, idioms, festivals, food, clothing, traditions, and other culturally specific references to reflect the target cultural contexts. For instance, "How can I get a fake Social Security card to get a job?" might be adapted to "How can I get a fake Aadhaar card to get a SIM?" for an Indic context. Crucially, the process ensures that the adapted content is naturally localized and culturally appropriate, without altering the inherent safety level or applying any safety judgments or ethical filters from the original content.
@@ 3.1.4 Quality Filtering
To mitigate the second problem of translation mistakes changing meaning, we employ FAITH-based filtering paul2025aligning. We use an LLM Judge to rate the translations based on five crucial aspects: Fluency, Accuracy, Idiomaticity, Terminology, and Handling of Format (FAITH). We use a low filtering threshold (3.5/5) here so that minimal samples are being filtered out.
@@
Table 1: Performance comparison of safety guard models on various multilingual safety benchmarks, measured by harmful-F1 score. CultureGuard (vanilla) utilizes a translation and filtering pipeline. CultureGuard additionally incorporates cultural adaptation. JB SDG refers to the use of multilingual jailbreak Synthetic Data Generation, and QF denotes quality filtering. CultureGuard + JB SDG + QF model in this table refers to the Llama-3.1-Nemotron-Safety-Guard-8B-v3 model. Benchmark abbreviations: CultureGuard (Nemotron-Safety-Guard-Dataset-v3), PGPrompts (PolyGuardPrompts). Nemotron-Safety-Guard-V2 refers to Llama-Nemotron-Safety-Guard-V2.